# V2V-LLM: Vehicle-to-Vehicle Cooperative Autonomous Driving with Multi-Modal Large Language Models

### v3

@@ 3.4 Evaluation Metrics
4 V2V-LLM
We also propose a competitive baseline model, V2V-LLM, for this LLM-based collaborative driving problem, as shown in Fig. 3. Our model is a Multi-Modal LLM (MLLM) that takes the individual perception features of every CAV as the vision input, a question as the language input, and generates an answer as the language output.
4.1 LiDAR-based Input Features
@@
(a) No fusion
(b) Early fusion
(c) Intermediate fusion [50, 51, 52]
Figure 4: Feature encoder diagrams of the baseline methods from different fusion approaches.
4.2 LiDAR-based Multi-Modal LLM
@@
Training: We use 8 NVIDIA A100-80GB GPUs to train our model. Our V2V-LLM uses LLaVA-v1.5-7b [25]’s Vicuna [7] as the LLM backbone. To train our model, we initialize it by loading the pre-trained LLaVA-v1.5-7b [25]’s checkpoint. We freeze the LLM and the point cloud feature encoder, and finetune the projector and the LoRA [14] parts of the model. During training, we use batch size 32. Adam optimizer is adopted for training with a starting learning rate $52e-$ 5 and a cosine learning rate scheduler with a 3% warm-up ratio. For all other training settings and hyperparameters, we use the same ones from LLaVA-v1.5-7b [25].
5 Experiment
5.1 Baseline Methods
We follow V2V4Real [53] and V2X-Real [48] to establish a benchmark for our proposed V2V-QA dataset with experiments on baseline methods using different fusion approaches: no fusion, early fusion, intermediate fusion, and our proposed baseline, LLM fusion (Fig. 3). The baseline methods also adopt the same projector and LLM architecture as in our V2V-LLM but with different point cloud feature encoders, as shown in Fig. 4. In some driving sequences of V2X-split that have point clouds from roadside infrastructures, we also include them as perception input in the same way as using CAVs’ point clouds.
@@ 5.1 Baseline Methods
Intermediate fusion: Prior work CoBEVT [50], V2X-ViT [51], and AttFuse [52] propose cooperative detection models that merge feature maps from multiple CAVs via attention mechanisms. Such approaches require less communication cost and can still achieve good performance. In our benchmark, we extract the features from those cooperative detection models as the input tokens to the LLM.
LLM fusion: We categorize our proposed V2V-LLM as a new type of fusion method, LLM fusion, which lets each CAV perform its individual 3D object detection to extract the scene-level feature maps and object-level feature vectors, and uses the Multi-Modal LLM to fuse the features from multiple CAVs. This approach is related to the traditional late fusion method that performs individual 3D object detection and aggregates the results by non-maximum suppression (NMS). Instead of applying NMS, our method adopts LLM to perform more tasks than just detection.
@@
Table 3: V2V-LLM’s testing performance in V2V-QA’s V2V-split and V2X-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline. More detailed performance evaluation can be seen in the supplementary material.
5.2 Experimental Results
5.2.1 Grounding
Our V2V-LLM and baseline methods’ performance on V2V-QA’s $3$ types of grounding questions can be seen in Tab. 3 for V2V-split and V2X-split, respectively. CoBEVT [50] is not included in V2X-split’s result because V2X-Real [48] does not release its CoBEVT [50] baseline model. In average, V2V-LLM  achieves similar performance in V2V-split and outperforms all other baseline methods in V2X-split. Such results indicate that our V2V-LLM has a promising capability of fusing perception features from multiple CAVs to answer grounding questions.
5.2.2 Notable Object Identification
Tab. 3 show the performance on the notable object identification task (Q4). Our proposed V2V-LLM outperforms all other methods in both V2V-split and V2X-split. Compared with the aforementioned grounding tasks, this notable object identification task requires more spatial understanding ability to identify the objects close to the planned future waypoints. For such a task, our V2V-LLM, which lets the Multi-Modal LLM perform both perception feature fusion and question answering, achieves the best results.
5.2.3 Planning
Tab. 3 show the performance of the planning task (Q5) for V2V-split and V2X-split, respectively. Our proposed V2V-LLM outperforms other methods in this safety-critical task to generate a future trajectory that aims to avoid potential collisions. More planning performance evaluation can be seen in the supplementary material.
5.2.4 Communication Cost and Scaling Analysis
In our centralized setting, each CAV sends one scene-level feature map ($\leq 0.2$MB), one set of individual object detection result parameters ($\leq 0.003$MB), one question ($\leq 0.0002$MB) to the LLM computing node and receives one answer ($\leq 0.0002$MB) at each timestep. If there are $N_{v}$ CAVs and each asks $N_{q}$ questions, the communication cost of each CAV is $(0.2+0.003+(0.0002+0.0002)N_{q})=(0.203+0.0004N_{q})$ MB, and the cost of the LLM is $(0.2+0.003+(0.0002+0.0002)N_{q})N_{v}=(0.203N_{v}+0.0004N_{q}N_{v})$ MB, as shown in Tab. 4. Note that each CAV only needs to send the same features to the LLM computing node once at each timestep because the LLM node can save and reuse them to answer multiple questions from the same or different CAVs at the same timestep. Detailed scaling analysis can be seen in the supplementary material.
@@ 5.2.4 Communication Cost and Scaling Analysis
Table 4: Communication cost (MB) and scaling analysis. $N_{v}$: number of CAVs. $N_{q}$: number of questions asked by each CAV at each timestep.
5.2.5 Summary
Overall, V2V-LLM achieves the best results in the notable object identification and planning tasks, which are critical in autonomous driving applications. V2V-LLM also achieves competitive results in the grounding tasks. In terms of communication costs, V2V-LLM only increases communication costs by $1.5\%$ in comparison to other intermediate fusion baseline methods. More detailed evaluation and analysis can be seen in the supplementary material.
@@ 5.2.5 Summary
Table 5: Ablation study in V2V-QA’s V2V-split testing set.
5.3 Ablation Study
Input Features: We experiment with variants of our V2V-LLM model that use either only the scene-level feature maps or only the object-level feature vectors as the visual input. The ablation results can be seen in Tab. 5. Both types of features contribute to final performance in all QA tasks. In general, the object-level-only model outperforms the scene-level-only model. This implies that the object-level features are easier for LLM to digest, which is consistent with the results observed in the previous work with the TOKEN model [39].
@@ 5.3 Ablation Study
Training from Scratch: Tab. 5 shows that training from scratch achieves worse performance, meaning that pre-training with LLaVA’s VQA tasks improves our V2V-LLM’s performance in V2V-QA. More detailed ablation results can be seen in the supplementary material.
@@
Figure 5: V2V-LLM’s grounding results on V2V-QA’s testing set. Magenta $\circ$: reference locations in questions. Yellow $+$: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 6: V2V-LLM’s notable object identification and planning results on V2V-QA’s testing set. For notable object identification, Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow $+$ and Cyan $\times$: model identification outputs corresponding to CAV_EGO and CAV_1, respectively. For planning, Green line: future trajectories in ground-truth answers. Yellow curve and Cyan curve: model planning outputs corresponding to CAV_EGO and CAV_1, respectively.
5.4 Qualitative Results
@@
Fig. 5 shows our V2V-LLM’s grounding results and the ground truth with visualization on V2V-QA’s testing set. We can observe that our V2V-LLM  is able to locate the objects given the provided reference information for each of the $3$ types of grounding questions. Fig. 6’s left part shows our V2V-LLM’s notable object identification results. V2V-LLM  demonstrate its capability of identifying multiple objects near the planned future trajectories specified in the questions for each CAV. Fig. 6’s right part shows V2V-LLM’s planning results. Our model is able to suggest future trajectories that avoid potential collisions with nearby objects. Overall, the outputs of our model closely align with the ground-truth answers across all question types, indicating its robustness in cooperative autonomous driving tasks.
6 Conclusion
@@
In this work, we expand the research scope of cooperative autonomous driving by integrating the use of Multi-Modal LLM-based methods, aimed at improving the safety of future autonomous driving systems. We propose a new problem setting and create a novel V2V-QA dataset and benchmark that includes grounding, notable object identification, and planning question-answering tasks designed for varieties of cooperative driving scenarios. We propose a baseline model V2V-LLM that fuses each CAV’s individual perception information and performs visual and language understanding to answer driving-related questions from any CAV. Our proposed V2V-LLM outperforms all other baselines adopted from state-of-the-art cooperative perception algorithms in the notable object identification and planning tasks, and achieves competitive performance in the grounding tasks. These experimental results indicate that V2V-LLM is promising as a unified multi-modal foundation model that can effectively perform perception and planning tasks for cooperative autonomous driving. We will publicly release our V2V-QA  dataset and code to facilitate the open-source research, and believe it will bring the cooperative driving research to the next stage.
7 Acknowledgement
The authors thank Boyi Li, Boris Ivanovic, and Marco Pavone for valuable discussions and comments.
Supplementary Material
8 Detailed Evaluation Results
@@
Tabs. 6 and 7 summarize the detailed evaluation results of our V2V-LLM and other baseline methods in V2V-QA’s V2V-split and V2X-split. In addition, Tabs. 8 and 9 show the detailed planning performance. For the grounding task, our V2V-LLM achieves competitive results in V2V-split and outperforms all other baseline methods in V2X-split. More importantly, for the notable object identification task and the planning task, our V2V-LLM outperforms all other baseline methods in both V2V-split and V2X-split.
@@
Table 6: V2V-LLM’s testing performance in V2V-QA’s V2V-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
@@
Table 7: V2V-LLM’s testing performance in V2V-QA’s V2X-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
@@
Table 8: V2V-LLM’s planning performance in V2V-QA’s V2V-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface, and the second-best results are in underline.
@@
Table 9: V2V-LLM’s planning performance in V2V-QA’s V2X-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface, and the second-best results are in underline.
9 Detailed Communication Cost and Scaling Analysis
@@
Alternatively, one can also consider a decentralized setting that deploys one LLM in each CAV. In this setting, each CAV receives the features from all other CAVs and does not need to send or receive any questions or answers. The communication cost of each CAV is $(0.2+0.003)(N_{v}-1)=0.203(N_{v}-1)$ MB. Tab. 10 summarizes the communication cost and scaling analysis in the aforementioned settings. There could be more different decentralized settings. Which setting works best in terms of communication costs is beyond the current focus of our work.
@@
Decentralized $0.203(N_{v}-1)$
Table 10: Communication cost (MB) and scaling analysis. $N_{v}$: number of CAVs. $N_{q}$: number of questions asked by each CAV at each timestep.
10 Planning Results with Temporal Inputs
In the main paper, all experiments use point clouds at a single frame from each CAV as the visual input to the models. In this section, we experiment with feeding visual features from $3$ consecutive frames, the current one and the previous two, as the visual input to the models. Tab. 11 shows the planning results of the new setting together with the original setting from the main paper. In general, using visual inputs from multiple frames improves planning performance.

### v4

@@ III-D Evaluation Metrics
IV V2V-LLM
We propose a competitive baseline model, V2V-LLM, for this LLM-based collaborative driving problem, as shown in Figure 3. Our model is a Multimodal LLM (MLLM) that takes the individual perception features of every CAV as the vision input, a question as the language input, and generates an answer as the language output.
IV-A LiDAR-based Input Features
@@
IV-B LiDAR-based Multimodal LLM
@@
Training: We use 8 NVIDIA A100-80GB GPUs to train our model. Our V2V-LLM uses LLaVA-v1.5-7b [16]’s Vicuna [4] as the LLM backbone. To train our model, we initialize it by loading the pre-trained LLaVA-v1.5-7b [16]’s checkpoint. We freeze the LLM and the point cloud feature encoder, and finetune the projector and the LoRA [10] parts of the model. During training, we use batch size 32. Adam optimizer is adopted for training with a starting learning rate 2e52e-5 and a cosine learning rate scheduler with a 3% warm-up ratio. For all other training settings and hyperparameters, we use the same ones from LLaVA-v1.5-7b [16].
V Experiment
V-A Baseline Methods
We follow V2V4Real [34] and V2X-Real [31] to establish a benchmark for our proposed V2V-QA dataset with experiments on baseline methods using different fusion approaches: no fusion, early fusion, intermediate fusion, and our proposed baseline, LLM fusion (3). All baseline methods also adopt the same projector and LLM architecture as in our V2V-LLM but with different point cloud feature encoders. In some driving sequences of V2X-split that have point clouds from roadside infrastructures, we include them as perception input as well in the same way as using CAVs’ point clouds.
@@ V-A Baseline Methods
Intermediate fusion: Prior work CoBEVT [33], V2X-ViT [35], and AttFuse [36] propose cooperative detection models that merge feature maps from multiple CAVs via attention mechanisms. Such approaches require less communication cost and can still achieve good performance. In our benchmark, we extract both of the scene-level and object-level features from those cooperative detection models as the input tokens to the LLM.
LLM fusion: We categorize our proposed V2V-LLM as a new type of fusion method, LLM fusion, which lets each CAV perform its individual 3D object detection to extract the scene-level feature maps and object-level feature vectors, and uses the Multimodal LLM to fuse the features from multiple CAVs. This approach is related to the traditional late fusion method that performs individual 3D object detection and aggregates the results by non-maximum suppression (NMS). Instead of applying NMS, our method adopts LLM to perform more tasks than just detection.
TABLE III: V2V-LLM’s testing performance in V2V-QA’s V2V-split and V2X-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
@@
V-B Quantitative Results
V-B1 Performance
Grounding: Our V2V-LLM and baseline methods’ performance on V2V-QA’s $3$ types of grounding questions can be seen in Table III for V2V-split and V2X-split, respectively. CoBEVT [33] is not included in V2X-split’s result because V2X-Real [31] does not release its CoBEVT [33] baseline model. In average, V2V-LLM  achieves similar performance in V2V-split and outperforms all other baseline methods in V2X-split. Such results indicate that our V2V-LLM has a promising capability of fusing perception features from multiple CAVs to answer grounding questions.
Notable Object Identification: Table III show the performance on the notable object identification task (Q4). Our proposed V2V-LLM outperforms all other methods in both V2V-split and V2X-split. Compared with the aforementioned grounding tasks, this notable object identification task requires more spatial understanding ability to identify the objects close to the planned future waypoints. For such a task, our V2V-LLM, which lets the Multimodal LLM perform both perception feature fusion and question answering, achieves the best results.
Planning: Table III show the performance of the planning task (Q5) for V2V-split and V2X-split, respectively. Our proposed V2V-LLM outperforms other methods in this safety-critical task to generate a future trajectory that aims to avoid potential collisions.
V-B2 Communication Cost and Scaling Analysis
In our centralized setting, each CAV sends one scene-level feature map ($\leq 0.2$MB), one set of individual object detection result parameters ($\leq 0.003$MB), one question ($\leq 0.0002$MB) to the LLM computing node and receives one answer ($\leq 0.0002$MB) at each timestep. If there are $N_{v}$ CAVs and each asks $N_{q}$ questions, the communication cost of each CAV is $(0.2+0.003+(0.0002+0.0002)N_{q})=(0.203+0.0004N_{q})$ MB, and the cost of the LLM is $(0.2+0.003+(0.0002+0.0002)N_{q})N_{v}=(0.203N_{v}+0.0004N_{q}N_{v})$ MB, as shown in Table IV. Note that each CAV only needs to send the same features to the LLM computing node once at each timestep because the LLM node can save and reuse them to answer multiple questions from the same or different CAVs at the same timestep.
TABLE IV: Communication cost (MB) and scaling analysis. $N_{v}$: number of CAVs. $N_{q}$: number of questions asked by each CAV at each timestep.
@@ V-B2 Communication Cost and Scaling Analysis
V-B3 Summary
In general, V2V-LLM achieves the best results in the notable object identification and planning tasks, which are critical in autonomous driving applications. V2V-LLM also achieves competitive results in the grounding tasks. In terms of communication costs, V2V-LLM only increases communication costs by $1.5\%$ in comparison to other intermediate fusion baseline methods.
V-C Comparison to Non-LLM baseline
To compare our V2V-LLM with non-LLM baseline in V2V-QA, for perception, we use the detection results from V2V4Real [34]’s best cooperative detection model CoBEVT [33] checkpoint (the same one as the feature extractor in our experiment) and evaluate them in our Q1 (grounding at a location). For planning, since there is no prior cooperative planning research on V2V4Real [34], we use the BEV features from the same CoBEVT [33] model checkpoint as input to initialize and train a BEV-planner [15], which is the best non-LLM planning baseline method in prior work OmniDrive [27]. Our V2V-LLM still outperforms the non-LLM cooperative perception and planning baseline, as shown in Table V.
TABLE V: Perception and planning performance comparison with non-LLM baseline method.
Method Q1 Q5
F1 $\uparrow$ L2avg (m) $\downarrow$ CRavg (%) $\downarrow$
CoBEVT [33] + BEV-planner [15] 65.7 5.82 11.59
V2V-LLM (ours) 70.0 4.99 3.00
V-D Robustness Assessment
We follow V2X-ViT [35] to experiment on the impact of latency and sensor noise on positional errors. Our model is robust to these factors, as shown in Tables VI and VII.
TABLE VI: Experiments in V2V-split with communication latency.
Latency (s) Q1 Q2 Q3 QGr Q4 Q5
F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ L2avg (m) $\downarrow$ CRavg (%) $\downarrow$
1.0 69.3 29.7 18.9 39.3 55.0 5.26 4.09
0.4 69.7 30.3 20.1 40.0 56.0 5.09 3.49
0.3 69.8 30.7 20.6 40.4 57.2 5.07 3.31
0.2 69.8 30.8 20.8 40.5 57.7 5.05 3.21
0.1 69.8 30.7 21.1 40.5 59.4 5.02 3.05
0 70.0 30.8 21.2 40.7 59.7 4.99 3.00
TABLE VII: Experiments in V2V-split with positional errors. STD: standard deviation.
STD (m) Q1 Q2 Q3 QGr Q4 Q5
F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ F1 $\uparrow$ L2avg (m) $\downarrow$ CRavg (%) $\downarrow$
1.0 69.8 29.9 21.7 40.5 57.2 5.21 3.86
0.4 69.8 30.9 21.5 40.7 59.2 5.03 3.27
0.3 69.8 30.7 21.0 40.5 59.1 5.00 3.20
0.2 69.8 30.9 21.0 40.6 60.0 4.99 3.10
0.1 69.8 30.8 21.3 40.6 59.8 4.98 3.05
0 70.0 30.8 21.2 40.7 59.7 4.99 3.00
V-E Ablation Study
Input Features: We experiment with variants of our V2V-LLM model that use either only the scene-level feature maps or only the object-level feature vectors as the visual input. The ablation results can be seen in Table VIII. Both types of features contribute to final performance in all QA tasks. In general, the object-level-only model outperforms the scene-level-only model. This implies that the object-level features are easier for LLM to digest, which is consistent with the results observed in the previous work with the TOKEN model [25].
TABLE VIII: Ablation study in V2V-split.
@@ V-E Ablation Study
Training from Scratch: Table VIII also shows that training from scratch achieves worse performance, meaning that pre-training with LLaVA’s VQA tasks improves our V2V-LLM’s performance in V2V-QA.
@@
Figure 4: V2V-LLM’s grounding results on V2V-QA’s testing set. Magenta $\circ$: reference locations in questions. Yellow ++: model output locations. Green $\circ$: ground-truth answers.
@@
Figure 5: V2V-LLM’s notable object identification and planning results on V2V-QA’s testing set. For notable object identification, Magenta curve: planned future trajectories in questions. Green $\circ$: ground-truth notable object locations. Yellow ++ and Cyan $\times$: model identification outputs corresponding to CAV_EGO and CAV_1, respectively. For planning, Green line: future trajectories in ground-truth answers. Yellow curve and Cyan curve: model planning outputs corresponding to CAV_EGO and CAV_1, respectively.
V-F Qualitative Results
@@
Figure 4 shows our V2V-LLM’s grounding results and the ground truth with visualization on V2V-QA’s testing set. We can observe that our V2V-LLM  is able to locate the objects given the provided reference information for each of the $3$ types of grounding questions: grounding at a reference location, grounding behind a reference object at a location, and grounding behind a reference object in a direction. Figure 5’s left part shows our V2V-LLM’s notable object identification results. V2V-LLM  demonstrate its capability of identifying multiple objects near the planned future trajectories specified in the questions for each CAV. Figure 5’s right part shows V2V-LLM’s planning results. Our model is able to suggest future trajectories that avoid potential collisions with nearby objects. Overall, the outputs of our model closely align with the ground-truth answers across all question types, indicating its robustness in cooperative autonomous driving tasks.
VI Conclusion
@@
In this work, we expand the research scope of cooperative autonomous driving by integrating the use of Multimodal LLM-based methods, aimed at improving the safety of future autonomous driving systems. We propose a new problem setting and create a novel V2V-QA dataset and benchmark that includes grounding, notable object identification, and planning question-answering tasks designed for varieties of cooperative driving scenarios. We propose a baseline model V2V-LLM that fuses each CAV’s individual perception information and performs visual and language understanding to answer driving-related questions from any CAV. Our proposed V2V-LLM outperforms other baselines adopted from state-of-the-art cooperative perception algorithms in the grounding, notable object identification, and planning. Our method also outperforms non-LLM baseline and is robust to communication latency and noise. These experimental results indicate that V2V-LLM is promising as a unified multimodal foundation model that can effectively perform perception and planning tasks for cooperative autonomous driving. We have publicly released our V2V-QA  dataset and V2V-LLM code to facilitate open-source research, and believe it will bring cooperative driving research to the next stage.
APPENDIX
VII Detailed Evaluation Results
@@
Tables IX and X summarize the detailed evaluation results of our V2V-LLM and other baseline methods in V2V-QA’s V2V-split and V2X-split. In addition, Tables XI and XII show the detailed planning performance. For the grounding task, our V2V-LLM achieves competitive results in V2V-split and outperforms all other baseline methods in V2X-split. More importantly, for the notable object identification task and the planning task, our V2V-LLM outperforms all other baseline methods in both V2V-split and V2X-split.
TABLE IX: V2V-LLM’s testing performance in V2V-QA’s V2V-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
@@
TABLE X: V2V-LLM’s testing performance in V2V-QA’s V2X-split in comparison with baseline methods. Q1: Grounding at a reference location. Q2: Grounding behind a reference object at a location. Q3: Grounding behind a reference object in a direction. QGr: Average of grounding (Q1, Q2, and Q3). Q4: Notable object identification. Q5: Planning. P: Precision. R: Recall. L2: L2 distance error. CR: Collision rate. Comm: Communication cost. In each column, the best results are in boldface, and the second-best results are in underline.
@@
TABLE XI: V2V-LLM’s planning performance in V2V-QA’s V2V-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface, and the second-best results are in underline.
@@
TABLE XII: V2V-LLM’s planning performance in V2V-QA’s V2X-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface, and the second-best results are in underline.
@@
VIII Detailed Communication Cost and Scaling Analysis
@@
Alternatively, one can also consider a decentralized setting that deploys one LLM in each CAV. In this setting, each CAV receives the features from all other CAVs and does not need to send or receive any questions or answers. The communication cost of each CAV is $(0.2+0.003)(N_{v}-1)=0.203(N_{v}-1)$ MB. Table XIII summarizes the communication cost and scaling analysis in the aforementioned settings. There could be more different decentralized settings. Which setting works best in terms of communication costs is beyond the current focus of our work.
TABLE XIII: Communication cost (MB) and scaling analysis. $N_{v}$: number of CAVs. $N_{q}$: number of questions asked by each CAV at each timestep.
@@
Decentralized $0.203(N_{v}-1)$ -
IX Planning Results with Temporal Inputs
In the main paper, all experiments use point clouds at a single frame from each CAV as the visual input to the models. In this section, we experiment with feeding visual features from $3$ consecutive frames, the current one and the previous two, as the visual input to the models. Table XIV shows the planning results of the new setting together with the original setting from the main paper. In general, using visual inputs from multiple frames improves planning performance.
TABLE XIV: V2V-LLM’s planning performance in V2V-QA’s V2V-split in comparison with baseline methods. L2: L2 distance error. CR: Collision rate. In each column, the best results are in boldface. and the second-best results are in underline.